# Large Language Model NLP Natural Language Processing

### v1

@@ 7 Conclusion
We presented the first study of the robustness of LLM-based deductive reasoning methods by introducing two types of perturbations: adversarial noise and counterfactual statements. These perturbations were used to examine the methodological aspects of LLM reasoners based on their format, syntax, and feedback mechanism for error recovery. While adversarial noise only affects autoformalisation approaches, counterfactual statements remain a significant challenge for all variants of the tested method. While feedback strategies may lead to fewer syntax errors in autoformalisation methods, the refined formalisations tend to be semantically incorrect, failing to increase accuracy. We call on future work to devise more advanced mechanisms for detecting, reporting, and incorporating semantic errors. We also anticipate generalizing the study in this paper to other logical deduction datasets.
Acknowledgements
@@ Appendix C Inference rules for counterfactual statements
Appendix D Implementation details
If available, the specific instruction-tuned variant has been used for the tested LLMs. Except for GPT 4o-mini, all models are provided from hugging-face and used with 4-bit quantization.
The Vampire theorem prover [^1] https://vprover.github.io processes the formal method’s input, generating proofs by refutation to check whether the query logically follows from the context.
Our FOL variant is adapted from the ANTLR grammars-v4 repository. [^2] https://github.com/antlr/grammars-v4/blob/5bb56f34bff3e0a43197d6e61ca89e2499e5c30b/fol/fol.g4 For each grammar in § 4.2, we generate a parser using ANTLR4333https://www.antlr.org/ to decouple the logical forms from the input requirements of the theorem prover.
For all error recovery strategies, the LLM prompt includes three in-context examples. When recovery is attempted, refinement is limited to three iterations. If the error persists, the random fallback strategy is applied. This setup allows us to systematically compare the robustness of these recovery methods in the autoformalisation process.
The generated warnings use three heuristics to identify semantic errors. First, checking for predicates and individuals only mentioned in the query to avoid an incomplete context. Second, ensuring that all predicates with the same identifier use the same amount of parameter (checking for narity). Finally, avoiding similar named predicates and individuals based on Levenshtein distance.
Appendix E Prompts
E.1 Direct prompt
@@ E.3 Autoformalistion prompt
(WalksInRain(john) $\lor$ $\neg$¬GetFit(john)) ::: John walks in the rain or he will not get fit.
@@ Conclusion:
GetWet(john) $\wedge$ $\neg$¬ExercisesALot(john) ::: John will get wet or he does not exercises a lot.
@@
$\neg$¬MissTrain(james) ::: James does not miss his train.

### v2

@@ 7 Conclusion
We presented the first study of the robustness of formal and informal LLM-based deductive reasoning methods by introducing two types of perturbations: adversarial noise and counterfactual statements. These perturbations were used to examine the methodological aspects of LLM reasoners based on their format, syntax, and feedback mechanism for error recovery. While adversarial noise only affects autoformalisation approaches, counterfactual statements remain a significant challenge for all variants of the tested method. While feedback strategies may lead to fewer syntax errors in autoformalisation methods, the refined formalisations tend to be semantically incorrect, failing to increase accuracy. We call on future work to devise more advanced mechanisms for detecting, reporting, and incorporating semantic errors. One possible mechanism to mitigate these issues could be the usage of models fine-tuned on error correction data. We also anticipate generalising the study in this paper to other logical deduction datasets and examining characteristics like the placement of adversarial noise and the effects of quantisation to strengthen the generalizability of the reported insights.
{ack}
@@ Appendix C Inference rules for counterfactual statements
Appendix D Prompts
D.1 Direct prompt
@@ D.3 Autoformalistion prompt
[…]
@@ Conclusion:
[…]
@@
[…]